OpenStudy (lena772):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

You can rearrange the given equation into point-slope form and then find the slope ... once you find the slope the parallel line should have the same number for slope and you can construct the equation using a new point-slope equation

OpenStudy (anonymous):

where point slope form is \[y - y_1 = m(x - x_1)\] which is a line of slope m that passes through point (x1, y1)

OpenStudy (lena772):

y--2=m(x-4) y+2=m(x-4)

OpenStudy (anonymous):

rearranging the given equation you should find y = 3(x + 1)

OpenStudy (anonymous):

So you find the slope of 3 here and apply it to the equation you found above